Two less than one-sixth of a certain number is 1 more than one-fourth that number. What is the number?
Answer by checkley71(8403)

X/6-2=X/4+1
X/6-X/4=1+2
(2X-3X)/12=3 NOW CROSS MULTIPLY.
-X=36
X=-36 ANSWER.
PROOF
-36/6-2=-36/4+1
-6-2=-9+1
-8=-8